OpenStudy (anonymous):

Find the average rate of change f(x)=x^3-3x+9 (a) from 5 to 9

OpenStudy (anonymous):

College Algebra

jimthompson5910 (jim_thompson5910):

Use the formula m = ( f(x2) - f(x1) )/( x2 - x1 ) where m = average rate of change x1 = first x value x2 = second x value f(x) is any function

jimthompson5910 (jim_thompson5910):

RONNCC you're thinking of the average of the function from x = 5 to x = 9 NOT the average rate of change from x = 5 to x = 9

jimthompson5910 (jim_thompson5910):

the two are completely different things

OpenStudy (anonymous):

Yea I tried plugging both in but still couldn't figure it out. ;(

jimthompson5910 (jim_thompson5910):

m = ( f(x2) - f(x1) )/( x2 - x1 ) m = ( f(9) - f(5) )/( 9 - 5 )

jimthompson5910 (jim_thompson5910):

m = ???

OpenStudy (anonymous):

Dont I need to know f first?

jimthompson5910 (jim_thompson5910):

f(x)=x^3-3x+9

jimthompson5910 (jim_thompson5910):

that's given

OpenStudy (anonymous):

I got all reals...

OpenStudy (anonymous):

@jim_thompson5910 nvm. it makes little sense why you would want to do that. your way is likely what is wanted.

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean by the average of the derivative

OpenStudy (anonymous):

I am confused now

jimthompson5910 (jim_thompson5910):

f(x)=x^3-3x+9 f(9)=(9)^3-3(9)+9 f(9) = ???

OpenStudy (anonymous):

so you're drawing a line from the beginning to the end, well how much is that line changing usually per unit 'x'? that's the average. so in this case (f(9)-f(5)) / (9-5) if you draw it out it may help

OpenStudy (anonymous):

711

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

f(9)=711

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

f(5) = ???

OpenStudy (anonymous):

119

OpenStudy (anonymous):

@jim_thompson5910 f(5)=119

jimthompson5910 (jim_thompson5910):

so put this all together to get m = ( f(x2) - f(x1) )/( x2 - x1 ) m = ( f(9) - f(5) )/( 9 - 5 ) m = ( 711 - 119 )/( 9 - 5 ) m = ???

OpenStudy (anonymous):

148

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

Sweet!